A neighborhood grocer sells a mix of chocolate and carob candy. The chocolate cost $2.70 a pound and the carob costs $2.55 a pound. If 20 pounds of the chocolate and 40 pounds of the carob candy are used, what is the cost per pound of the mixture?
Answer by checkley71(8403)

2.70*20+2.55*40=60X OR 54.00+102.00=60X OR 156.00=60X OR 156.00/60=X OR X=2.60